OpenStudy (anonymous):

hey anyone knows why can`t get the exact answer of these questions coz when i solved it it turns my answer is 119/12 for problem #1 and 15/22 for problem #2. these are the problems 1.) 3 1/2 x 2 5/6=n choices: a.6 2/9 b. 9 4/9 c. 5 6/9 d. 6 5/18 and 2.) 2 3/3 divide 4 2/5 choices: a.19/24 b. 11/12 c. 3/4 d. 5/8 please show me the solution on how to get those anwers.

OpenStudy (anonymous):

1. 3 1/2 x 2 5/6 (turn the mixed numbers into improper fractions) 7/2 x 17/6 = 119/12 = 9 11/12 Are you sure you wrote the answer choices correctly ? Or even the problem ? 2. 2 3/3 divided by 4 2/5 (change to improper fractions) 9/3 divided by 22/5 9/3 x 5/22 = 45/66 = 15/22 I don't know about these problems....either you wrote them down wrong, or the book is wrong.

OpenStudy (anonymous):

that`s the problem but the choices aren`t the same as my answer.

OpenStudy (anonymous):

They are not the same as my answers either. I don't know what to say :/

OpenStudy (anonymous):

Do you take school on a computer ?

OpenStudy (anonymous):

maybe those problems are mathematical error? coz our answer aren`t in the choices..but anyway thanks for the effort and time for solving those problems.

OpenStudy (anonymous):

no problem....at least you know you are not doing them wrong....you are not crazy..lol :)

OpenStudy (anonymous):

hahaha..i`ll try to ask my instructor for these problems.thanks again.